OpenStudy (anonymous):

Determine all of the roots of the given equation: x^3-2x^2-3x+10=0

OpenStudy (anonymous):

Complex, too? @mmbuckaroos?

OpenStudy (anonymous):

x=-2 x=2-i x=2+i

OpenStudy (anonymous):

Could you explaint o me how you got those so I can try my other problems please?

OpenStudy (anonymous):

Sorry, I'm a little low on time... if you bump your question, i'm sure someone else can help you. sorry!

OpenStudy (anonymous):

ok thanks

OpenStudy (mertsj):

The possible rational roots are the factors of the constant term which is 10 divided by the factors of the coefficient of the term with the highest exponent which is the x^3 term. It's coefficient is 1. So the possible rational roots are: \[\frac{\pm10}{1}, \frac{\pm5}{1}, \pm\frac{2}{1}, \frac{\pm1}{1}\]

OpenStudy (mertsj):

Use synthetic division to see if any of those are actually roots.

OpenStudy (mertsj):

Since we have reason to believe that -2 is a root, let's try it.

OpenStudy (mertsj):